# Benjamin Carleton

### v6

@@ Abstract
Electoral control types are ways of trying to change the outcome of elections by altering aspects of their composition and structure [BTT92]. We say two compatible (i.e., having the same input types) control types that are about the same election system $\mathcal{E}$ form a collapsing pair if for every possible input (which typically consists of a candidate set, a vote set, a focus candidate, and sometimes other parameters related to the nature of the attempted alteration), either both or neither of the attempted attacks can be successfully carried out (see the Preliminaries for a more formal definition) [HHM20]. For each of the seven general (i.e., holding for all election systems) electoral control type collapsing pairs found by Hemaspaandra, Hemaspaandra, and Menton [HHM20] and for each of the additional electoral control type collapsing pairs of Carleton et al. [${}^{+}$22] for veto and approval (and many other election systems in light of that paper’s Theorems 3.6 and 3.9), both members of the collapsing pair have the same complexity since as sets they are the same set. However, having the same complexity (as sets) is not enough to guarantee that as search problems they have the same complexity. In this paper, we explore the relationships between the search versions of collapsing pairs. For each of the collapsing pairs of Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [${}^{+}$22], we prove that the pair’s members’ search-version complexities are polynomially related (given access, for cases when the winner problem itself is not in polynomial time, to an oracle for the winner problem). Beyond that, we give efficient reductions that from a solution to one compute a solution to the other. For the concrete systems plurality, veto, and approval, we completely determine which of their (due to our results) polynomially-related collapsing search-problem pairs are polynomial-time computable and which are NP-hard.
@@ 1 Introduction
Algorithms for problems on elections are extremely important combinatorial algorithms, where attackers generally have available an exponential number of potential actions. Despite the combinatorially explosive number of potential actions available, in many cases polynomial-time algorithms can be developed. Thus the study of manipulative actions against elections is a beautiful showcase of an area with real-world importance, where the outcome of the clash between finding efficient algorithms for combinatorially explosive problems, and proving the impossibility (if $\mathrm{P}\neq\mathrm{NP}$) of finding such algorithms, is of vivid interest.
@@
Hemaspaandra, Hemaspaandra, and Menton [HHM20] noted that, surprisingly, seven pairs among the 44 (relatively) “standard” control types collapse: For each election system (i.e., each mapping from candidates and votes to a winner set among the candidates) $\mathcal{E}$ whose vote type is linear orders, for each of those seven pairs of types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ it holds that for each input there is a successful action on that input under the $\mathcal{T}_{1}$ control type if and only if there is a successful action on that input under the $\mathcal{T}_{2}$ control type. Viewed as sets, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are identical! Carleton et al. [${}^{+}$22] noted that Hemaspaandra, Hemaspaandra, and Menton’s proof tacitly in fact established that those seven collapses hold for all election systems regardless of their vote type.
Carleton et al. [${}^{+}$22] showed that those seven pairs are the only collapsing pairs among the 44 standard control types, at least if one wants the given collapse to hold for every election system. However, Carleton et al. [${}^{+}$22] discovered some additional collapses that hold specifically for veto or specifically for approval voting, and also found some additional collapses that hold for all election systems that satisfy certain axiomatic properties.
For each election system to which a collapse applies, the complexity of the two collapsing control types is the same (since due to collapsing, the two control types yield the exact same set). However, Carleton et al. [${}^{+}$22] raised (even in its initial technical report version, which was the main motivation for the present paper) as a challenge the issue of whether the search complexity of the collapsing pairs is also the same. That is the issue that is the focus of the present paper.
@@
This paper explores the issue of whether the collapsing (as decision problems) pairs of Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [${}^{+}$22] have the same search complexity. We prove that in every case the answer is Yes (if given access, for the case where the election’s winner problem itself is not even in $\mathrm{P}$, to an oracle for the election’s winner problem), and indeed we show that in each case a solution for one can (again, given access to the winner problem) be polynomial-time transformed into a solution for the other. Thus the complexities are polynomially related (given access to the election system’s winner problem). Additionally, for the concrete cases of plurality’s, veto’s, and approval’s collapsing pairs, we explore whether those polynomially equivalent search complexities are clearly “polynomial time,” or are $\mathrm{NP}$-hard, and we resolve every such case.
Why is this important? In reality, one typically—e.g., if one is a campaign manager—wants not merely to efficiently compute whether there exists some action that will make one’s candidate win, but rather one wants to get one’s hands, efficiently, on an actual such successful action. Unfortunately, the literature’s existing collapsing control type pairs are each about and proved for the “exists” case. In contrast, this paper’s results are establishing that the literature’s existing collapsing control type pairs even have the property that—given access to the winner problem for the election system in question [^1] Although for the three concrete election systems we cover as examples that is not even needed as their winner problems are each in polynomial time.—for both members of the pair the “getting one’s hands on a successful action when one exists” issue is of the same complexity for both. We also show that one usually can efficiently use access to solutions for one to get solutions for the other.
@@ 2.1 Elections and Election Systems
An election consists of a finite candidate set $C$ and a finite collection of votes $V$ over the candidates in $C$. The “type” of the votes depends on the election system one is considering. Most typically, each vote is a linear ordering—a complete, transitive, asymmetric binary relation—over the candidates (e.g., $3>1>2$). Another common vote type (called an approval vector) is that a vote is bit-vector of length $\|C\|$, with each bit typically indicating approval (1) or disapproval (0) of a candidate. No two candidates can have the same name. [^2] By allowing candidates to have names, we are following the model of the papers that this paper is most closely related to [HHM20, ${}^{+}$22]. It is possible that election systems in this model may exploit candidate names in complicated ways. However, adopting this model in fact makes our results stronger than they would be if our model for example required election systems to be (candidate-)neutral and/or to have the candidate identities in $(C,V)$ always be $1,2,\dots,\|C\|$. Also, studies on control are incompatible with assuming that candidate names are always $1,2,\dots,\|C\|$, since many control types, including all partition-based ones, change the candidate set, thus resulting in elections that do not satisfy that condition; that issue has been implicit ever since the seminal work of Bartholdi, Tovey, and Trick [BTT92] that initiated the study of control attacks on elections.
@@ 2.2 Control Types and Collapses
Definition 2.1 and the text following it define the 24 partition-based control types (for each election system $\mathcal{E}$). For uniformity, we take the definition essentially verbatim from the papers we are most closely related to [HHM20, ${}^{+}$22], which themselves were drawing on the line of earlier papers—starting with Bartholdi, Tovey, and Trick [BTT92]—that developed the current set of control notions (for more history and citations, see [FR16, HHM20, ${}^{+}$22]). One must be clear as to the handling of candidates who are tied winners in first-round elections. The two tie-handling rules typically studied as to partition-based control types are ties-eliminate ($\mathrm{TE}$)—in which a candidate must uniquely win its subelection to proceed to the next round—and ties-promote ($\mathrm{TP}$)—in which all winners of a subelection proceed to the next round. In Definition 2.1 and later in the paper, we will at times have elections whose candidate set is $C^{\prime}$ but whose votes, due to candidate partitioning and/or first-round candidate eliminations, are over a set $C\supseteq C^{\prime}$. As is standard in the literature, in such cases we always take this to mean that the votes are each masked down to just the candidates in $C^{\prime}$.
Definition 2.1 (see [HHM20, ${}^{+}$22] and the references/history therein)
@@
In each of the six control types defined above we can replace “is a winner” with “is a unique winner” to denote constructive control in the unique-winner ($\mathrm{UW}$) model. The resulting control types are appended with “-$\mathrm{UW}$” rather than “-$\mathrm{NUW}$.” $\mathrm{NUW}$ denotes the so-called nonunique-winner model, aka, the cowinner model, in which merely being an overall winner, whether tied or not, is the goal. This completes the definition of the 12 standard constructive partition-control types (for each election system; in text discussions, we will often omit the “for each election system,” and so may speak of control of a given type generically). The standard 12 additional destructive partition-control types differ only in that instead of seeking to make a focus candidate a winner or unique winner, the goal is to prevent the focus candidate from being a winner or unique winner. To denote those, we replace the “$\mathrm{CC}$” with a “$\mathrm{DC}$,” e.g., the destructive control by partition of voters problem for $\mathcal{E}$ using the TE handling rule and the UW model is denoted $\mathcal{E}$-DC-PV-TE-UW. Thus the 24 standard types of partition-based control are defined. As is standard, each control type is quietly defining—in some sense is—a set, namely, the set of all inputs on which the given control attack can be successfully carried out. Thus for example $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ is a set. We will immediately use this view in the next sentence.
@@
For any two control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ that are compatible (i.e., have the same collection of input fields), Carleton et al. [${}^{+}$22] say that $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ collapse if $\mathcal{T}_{1}=\mathcal{T}_{2}$. If $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are both about the same election system $\mathcal{E}$ (i.e., each has as its prefix “$\mathcal{E}$-”) and $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are compatible, we will say that control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are $\mathcal{E}$-matched. If control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are both about the same election system $\mathcal{E}$ and collapse (and so also are compatible), we say that they are a collapsing pair. Since all 24 control types just defined are mutually compatible, and none of the other $44-24=20$ standard control types are known to be involved in any of the collapses found in Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [${}^{+}$22], we will not further mention compatibility in this paper. (See [${}^{+}$22] for the standard definitions of the 20 control types not defined here.)

### v7

# Search Versus Search for Collapsing Electoral Control Types

@@ Abstract
Electoral control types are ways of trying to change the outcome of elections by altering aspects of their composition and structure [BTT92]. We say two compatible (i.e., having the same input types) control types that are about the same election system $\mathcal{E}$ form a collapsing pair if for every possible input (which typically consists of a candidate set, a vote set, a focus candidate, and sometimes other parameters related to the nature of the attempted alteration), either both or neither of the attempted attacks can be successfully carried out (see the Preliminaries for a more formal definition) [HHM20]. For each of the seven general (i.e., holding for all election systems) electoral control type collapsing pairs found by Hemaspaandra, Hemaspaandra, and Menton [HHM20] and for each of the additional electoral control type collapsing pairs of Carleton et al. [CCH+24] for veto and approval (and many other election systems in light of that paper’s Theorems 3.6 and 3.9), both members of the collapsing pair have the same complexity since as sets they are the same set. However, having the same complexity (as sets) is not enough to guarantee that as search problems they have the same complexity. In this paper, we explore the relationships between the search versions of collapsing pairs. For each of the collapsing pairs of Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [CCH+24], we prove that the pair’s members’ search-version complexities are polynomially related (given access, for cases when the winner problem itself is not in polynomial time, to an oracle for the winner problem). Beyond that, we give efficient reductions that from a solution to one compute a solution to the other. For the concrete systems plurality, veto, and approval, we completely determine which of their (due to our results) polynomially-related collapsing search-problem pairs are polynomial-time computable and which are NP-hard.
@@ 1 Introduction
Algorithms for problems on elections are an important class of combinatorial algorithms, in which an election’s attacker generally has available an exponential number of potential actions. Despite the combinatorially explosive number of potential actions available, in many cases polynomial-time algorithms can be developed. Thus the study of manipulative actions against elections is a beautiful showcase of an area with real-world importance, where the outcome of the clash between finding efficient algorithms for combinatorially explosive problems, and proving the impossibility (if $\mathrm{P}\neq\mathrm{NP}$) of finding such algorithms, is of vivid interest.
@@
Hemaspaandra, Hemaspaandra, and Menton [HHM20] noted that, surprisingly, seven pairs among the 44 (relatively) “standard” control types collapse: For each election system (i.e., each mapping from candidates and votes to a winner set among the candidates) $\mathcal{E}$ whose vote type is linear orders, for each of those seven pairs of types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ it holds that for each input there is a successful action on that input under the $\mathcal{T}_{1}$ control type if and only if there is a successful action on that input under the $\mathcal{T}_{2}$ control type. Viewed as sets, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are identical! Carleton et al. [CCH+24] noted that Hemaspaandra, Hemaspaandra, and Menton’s proof tacitly in fact established that those seven collapses hold for all election systems regardless of their vote type.
Carleton et al. [CCH+24] showed that those seven pairs are the only collapsing pairs among the 44 standard control types if one wants the given collapse to hold for every election system. However, that paper discovered some additional collapses that hold specifically for veto or for approval voting, and also found new collapses that hold for all election systems that satisfy certain axiomatic properties. [^1] Two of those axiomatic theorems had incorrect proofs in the Carleton et al. conference version [CCH+23b] and/or its associated arXiv.org full technical report [CCH+22], but the journal version of that paper [CCH+24], and the arXiv.org full technical report’s May 2024 revisions, change the axiomatic assumptions and then correctly prove the thus modified results. The conference version was using the flawed results to prove things about approval voting, and fortunately the journal version, and the arXiv.org technical report’s May 2024 revisions, show that those consequences can still be obtained using the modified results.
For each election system to which a collapse applies, the complexity of the two collapsing control types is the same (since, due to collapsing, the two control types yield the exact same set). However, Carleton et al. [CCH+24] raised (even in its initial technical report version, which was the main motivation for the present paper) as a challenge the issue of whether the search complexity of the collapsing pairs is also the same. That is the issue that is the focus of the present paper.
@@
This paper explores the issue of whether the collapsing (as decision problems) pairs of Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [CCH+24] have the same search complexity. We prove that in every case the answer is Yes (if given access, for the case where the election’s winner problem itself is not even in $\mathrm{P}$, to an oracle for the election’s winner problem), and indeed we show that in each case a solution for one can (again, given access to the winner problem) be polynomial-time transformed into a solution for the other. Thus the complexities are polynomially related (given access to the election system’s winner problem). Additionally, for the concrete cases of plurality’s, veto’s, and approval’s collapsing pairs, we explore whether those polynomially equivalent search complexities are clearly “polynomial time,” or are $\mathrm{NP}$-hard, and we resolve every such case.
Why is this important? In reality, one typically—e.g., if one is a campaign manager or a corrupt vote collector/counter—wants not merely to efficiently compute whether there exists some action that will make one’s candidate win, but rather one wants to get one’s hands, efficiently, on an actual such successful action. Unfortunately, the literature’s currently known results on collapsing control type pairs are each about the “exists” case. In contrast, this paper’s results are establishing that the literature’s existing collapsing control type pairs even have the property that—given access to the winner problem for the election system in question [^2] Although for the three concrete election systems we cover as examples that is not even needed as their winner problems are each in polynomial time.—for both members of the pair the “getting one’s hands on a successful action when one exists” issue is of the same complexity for both. We also show that one usually can efficiently use access to solutions for one to get solutions for the other.
@@ 2.1 Elections and Election Systems
An election consists of a finite candidate set $C$ and a finite collection of votes $V$ over the candidates in $C$. The “type” of the votes depends on the election system that one is considering. Most typically, each vote is a linear ordering—a complete, transitive, asymmetric binary relation—over the candidates (e.g., $3>1>2$). Another common vote type (called an approval vector) is that a vote is bit vector of length $\|C\|$, with each bit typically indicating approval (1) or disapproval (0) of a candidate. No two candidates can have the same name. [^3] By allowing candidates to have names, we are following the model of the papers that this paper is most closely related to [HHM20, CCH+24]. It is possible that election systems in this model may exploit candidate names in complicated ways. However, adopting this model in fact makes our results stronger than they would be if our model, for example, required election systems to be (candidate-)neutral and/or to have the candidate identities in $(C,V)$ always be $1,2,\dots,\|C\|$. Also, studies on control are incompatible with assuming that candidate names are always $1,2,\dots,\|C\|$, since many control types, including all partition-based ones, change the candidate set, thus resulting in elections that do not satisfy that condition; that issue has been implicit ever since the seminal work of Bartholdi, Tovey, and Trick [BTT92] that initiated the study of control attacks on elections.
@@ 2.2 Control Types and Collapses
Definition 2.1 and the text following it define the 24 partition-based control types (for each election system $\mathcal{E}$). For uniformity, we take the definition essentially verbatim from the papers we are most closely related to [HHM20, CCH+24], which themselves were drawing on the line of earlier papers—starting with Bartholdi, Tovey, and Trick [BTT92]—that developed the current set of control notions (for more history and citations, see [FR16, HHM20, CCH+24]). One must be clear as to the handling of candidates who are tied winners in first-round elections. The two tie-handling rules typically studied as to partition-based control types are ties-eliminate ($\mathrm{TE}$)—in which a candidate must uniquely win its subelection to proceed to the next round—and ties-promote ($\mathrm{TP}$)—in which all winners of a subelection proceed to the next round. In Definition 2.1 and later in the paper, we will at times speak of elections whose candidate set is $C^{\prime}$ but whose votes, due for example to candidate partitioning and/or first-round candidate eliminations, are over a set $C\supseteq C^{\prime}$. For example, for the $(C_{1},V)$ of part 2 of Definition 2.1, the $C^{\prime}$ is $C_{1}$ and the votes in $V$ are over $C$. As is standard in the literature (e.g., this was used implicitly in the seminal control paper [BTT92] and many other control papers since, see also [CCH+24, Footnote 4] where it is explicit), in such cases we always take this to mean that the votes are each masked down to just the candidates in $C^{\prime}$.
Definition 2.1 (see [HHM20, CCH+24] and the references/history therein)
@@
In each of the six control types defined above, we can replace “is a winner” with “is a unique winner” to denote constructive control in the unique-winner ($\mathrm{UW}$) model. The resulting control types are appended with $\mathrm{UW}$” rather than $\mathrm{NUW}$.” $\mathrm{NUW}$ denotes the so-called nonunique-winner model, aka, the cowinner model, in which merely being an overall winner, whether tied or not, is the goal. This completes the definition of the 12 standard constructive partition-control types (for each election system; in text discussions, we will often omit the “for each election system,” and so may speak of control of a given type generically). The standard 12 additional destructive partition-control types differ only in that instead of seeking to make a focus candidate a winner or unique winner, the goal is to prevent the focus candidate from being a winner or unique winner. To denote those, we replace the $\mathrm{CC}$” with a $\mathrm{DC}$,” e.g., the destructive control by partition of voters problem for $\mathcal{E}$ using the TE handling rule and the UW model is denoted $\mathcal{E}$-DC-PV-TE-UW. Thus the 24 standard types of partition-based control are defined. As is standard, each control type is quietly defining—in some sense is—a set, namely, the set of all inputs on which the given control attack can be successfully carried out. Thus, for example, $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ is a set. We will immediately use this view in the next sentence.
@@
For any two control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ that are compatible (i.e., have the same collection of input fields), Carleton et al. [CCH+24] say that $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ collapse if $\mathcal{T}_{1}=\mathcal{T}_{2}$. If $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are both about the same election system $\mathcal{E}$ (i.e., each has as its prefix $\mathcal{E}$-”) and $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are compatible, we will say that control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are $\mathcal{E}$-matched. If control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are both about the same election system $\mathcal{E}$ and collapse (and so also are compatible), we say that they are a collapsing pair. Since all 24 control types just defined are mutually compatible, and none of the other 4424=2044-24=20 standard control types are known to be involved in any of the collapses found in Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [CCH+24], we will not further mention compatibility in this paper. (See [CCH+24] for the standard definitions of the 20 control types not defined here.)